OpenStudy (anonymous):

A 2-lb bag of brown rice costs seven dollars. How much would you expect to pay for a ten-lb bag of brown rice, assuming there is no discount for buying a larger bag? fourteen dollars fifteen dollars twenty-eight dollars and fifty-seven cents thirty-five dollars

OpenStudy (zarkon):

\[\frac{2}{7}=\frac{10}{x}\] solve for x

OpenStudy (saifoo.khan):

2x = 70 x = 70/2 x = 35

OpenStudy (anonymous):

hey can u guys do this question?